OpenStudy (anonymous):

what's the formula for finding (pi) in terms of the circumference and radius?

OpenStudy (anonymous):

In circles the AREA is equal to 3.14 (p) times the radius (r) to the power of 2. Thus the formula looks like: A= pr2.

OpenStudy (zepp):

I think he's asking for the circumference, 2pir, not area :s

OpenStudy (anonymous):

π = C + 2r is that it?

OpenStudy (anonymous):

Oh. xD I fail because that's how I learned it. Nevermind then. @Student890 You got it right. That's what I meant. But i learned both. Dx Fail.

OpenStudy (zepp):

pi = c/2r since C = 2pir

OpenStudy (anonymous):

Oh don't worry about it! I'm not sure if it's that one or π = C divided by two r

OpenStudy (zepp):

It's π = C divided by two r